OpenStudy (anonymous):

6x - 2y=4 y + 2=3x.....I need help showing that this is a dependent equations.

OpenStudy (anonymous):

Ok first we need to solve for x or y in one equation so lets start off by doing this making y+2=3 into y=3x-2

OpenStudy (anonymous):

Now we know what y equals so we then plug in y into 6x-2y=4

OpenStudy (anonymous):

Are you getting it so far?

OpenStudy (anonymous):

so i plug y=3x-2 after the 2 in 6x-2y=4?

OpenStudy (anonymous):

yes :) so it will be 2(3x-2)

OpenStudy (anonymous):

Because y =3x-2 you can plug in that after the 2

OpenStudy (anonymous):

Now solve for x

OpenStudy (anonymous):

and tell me what you get

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

ok lemme check your awnser

OpenStudy (anonymous):

Yes that is correct :D

OpenStudy (anonymous):

So this proves that one equation is dependant on the other equations set of values

OpenStudy (anonymous):

then i plug the 0 for x in 6x-2y=4 to solve for y?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so how would it look like when I graph it?

OpenStudy (anonymous):

0=3x-2 2=3x x=2/3

OpenStudy (anonymous):

Lemme refresh my memory a bit sorry

OpenStudy (anonymous):

Actually we are sposed to solve both equations for Y, silly me. Ok so we know y=3x-2 but also we need to solve for y in 6x-2y=4 So we change 6x-2y=4 to -2y=4-6x Then we divide both sides by -2 to get y=3x-2 So since they both are the same line you just graph 3x-2

OpenStudy (anonymous):

Sorry about that I was in the wrong mindset

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

So it should look something like this 

OpenStudy (anonymous):

Because both equations produce the same line and are consistant so therefore they are dependant of eachother Its a phrase you can use if you are instructed to explain your awnser

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

Not a problem :)

OpenStudy (anonymous):

You can add me as a fan if you wish for my help whenever im online

OpenStudy (anonymous):

Ill be glad to help you again

OpenStudy (anonymous):

Thanks again for your help